OpenStudy (anonymous):

Factor completely: x^2 − 3x − 10

OpenStudy (anonymous):

(x-5)(x-2)

OpenStudy (anonymous):

is that equal to zero

OpenStudy (anonymous):

()x-5)(x+2)=0

OpenStudy (anonymous):

Factor the equation to (x - 5)(x + 2) = 0 Set both factors equal to 0 x - 5 = 0 x + 2 = 0 Solve for x and the answer is x = 5 x = -2